Table-tennis ball has a mass 10 g and a speed of 90 m/s. If speed can be measured within an accuracy of 4% what will be the uncertainty in speed and position?

1
Expert's answer
2020-07-27T05:32:41-0400

Given, mass of ball = 10g = 10* 10-3 kg

speed = 90m/s

accuracy = 4%

so uncertainty in speed = 4% of 90 = 4*90/100 = 3.6 m/s


so using uncertainty formula


"\Delta x * \Delta v = h/4\pi m"

"\Delta x= 6.626*10^{-34}/(4*3.14*10^{-2}*3.6)"

"\Delta x = 1.46 * 10 ^{-33}" m